OpenStudy (anonymous):

If cos/theta= -5/13 and sin/theta < 0, what is tan/theta ? a) 12/5 b) 13/5 c) -12/5 d) 194/5

OpenStudy (anonymous):

a)12/5 sin(theta)=-12/13 so tan(theta)=sin/cos=12/5